1. 7hot dogs and 4 hamburgers cost $15.00 how much does 1 Hot dog cost?
2. 4 hotdogs and 7 hamburgers $18.00 how much does 1 hamburger cost?

Answer by stanbon(75887)

1. 7hot dogs and 4 hamburgers cost $15.00 how much does 1 Hot dog cost?
2. 4 hotdogs and 7 hamburgers $18.00 how much does 1 hamburger cost?
---------------------
Equations:
7t + 4h = 15
4t + 7h = 18
------
Modify for elimination:
28t + 16h = 60
28t + 49h = 126
-----
33h = 66
h = $2.00 (cost of one hamburger)
-------
Solve for "t":
4t + 7h = 18
4t + 7*2 = 18
4t = 4
t = $1.00 (cost of one hot dog)
================================
Cheers,
Stan H.